Find two distinct points on the curve y=x3+x2-x+1 that have parallel tangent lines


1
Expert's answer
2021-10-20T14:22:51-0400

Given curve is "y=x^3+x^2-x+1."

"\therefore \frac{dy}{dx}=3x^2+2x-1"

Let the two points which have parallel tangents lines be "(x_1,y_1)" and "(x_2,y_2)."

Therefore, slope of both tangents must be equal.

"3{x_1}^2+2{x_1}-1=3{x_2}^2+2{x_2}-1\\
\Rightarrow 3({x_1}^2-{x_2}^2)+2({x_1}-{x_2})=0\\
\Rightarrow ({x_1}-{x_2})(3({x_1}+{x_2})+2)=0\\
\therefore x_1=x_2 \ or \ x_1+x_2=-\frac{2}{3}"

But, we know that "x_1\neq x_2" because than both the points will be same.

So, "\ x_1+x_2=-\frac{2}{3}"

Hence, we can conclude that every two points whose sum of abscissa is "-\frac{2}{3}" will satisfy this condition.